# Explorations of Self-Repair in Language Models

### v1

@@ Abstract
Prior interpretability research studying narrow distributions has preliminarily identified self-repair, a phenomena where if components in large language models are ablated, later components will change their behavior to compensate. Our work builds off this past literature, demonstrating that self-repair exists on a variety of models families and sizes when ablating individual attention heads on the full training distribution. We further show that on the full training distribution self-repair is imperfect, as the original direct effect of the head is not fully restored, and noisy, since the degree of self-repair varies significantly across different prompts (sometimes overcorrecting beyond the original effect). We highlight two different mechanisms that contribute to self-repair, including changes in the final LayerNorm scaling factor (which can repair up to 30% of the direct effect) and sparse sets of neurons implementing Anti-Erasure. We additionally discuss the implications of these results for interpretability practitioners and close with a more speculative discussion on the mystery of why self-repair occurs in these models at all, highlighting evidence for the Iterative Inference hypothesis in language models, a framework that predicts self-repair.
@@ 1 Introduction
We strengthen this prior work by investigating self-repair across the whole pretraining distribution, by focusing on individual attention heads (a smaller change, and thus more surprising to see repaired), and by investigating the mechanisms behind self-repair on the whole distribution. Our key findings are:
@@
    Up to 30% of the self-repairing of direct effects can be attributed to just the effect of ablations on the LayerNorm normalization factor. This is even more pronounced when using zero ablations.
@@
    MLP Erasure, a mechanism in which MLP layers ’erase’ outputs from earlier model components and is known to explain self-repair (McGrath et al., 2023), is powered by a sparse set of Erasure neurons. The exact neurons vary between prompts.
We end with some discussion on the implications of these results for interpretability research and present preliminary evidence for the Iterative Inference hypothesis, a speculative framework that predicts self-repair as a side effect. All of our code for the experiments used in this paper is provided at https://github.com/starship006/backup_research.
@@ 2 Self-Repair on the Full Distribution Exists, but is Incomplete and Noisy
In this section, we confirm that across models, individual attention heads are ’self-repaired’ on a general pretraining distribution. We detail how we measure self-repair, demonstrate how individual attention heads are imperfectly self-repaired across the entire pretraining distribution, and highlight how it is noisy.
2.1 Measuring Self-Repair
We first detail how we define direct effect and self-repair. Our methodology aligns with previous research into self-repair (Wang et al., 2023; McGrath et al., 2023), although not mirroring it identically. Individual model components can be thought of as having a direct effect: the effect on the logits, not mediated by any subsequent component (McGrath et al., 2023). We measure self-repair by comparing the change in the direct effect $\Delta\text{DE}_{\text{head}}$ of individual attention heads with the logit difference $\Delta\text{logit}$ of the correct next token once the attention head is ablated.
@@ 2.1 Measuring Self-Repair
Across each token, we measure a direct effect $DE_{\text{head}}$ of individual attention heads. Model components add their outputs into the residual stream. As such, the residual stream can be decomposed into the output of each model component (including individual attention heads). Because the residual stream (mostly) linearly maps into the logits, there exists an associated direction in the residual stream that maps to the correct next token. We can measure how much each attention head contributed to it (Elhage et al., 2021): this is the direct effect $DE_{\text{head}}$ of an attention head. Ideally, if something has a high direct effect, it was likely ’important’ for predicting the next token.
We contrast the direct effect with ablation-based metrics: in particular, we use resample ablations, a common technique that replaces the output of the same head on a different set of pretraining tokens (Chan et al., 2022). We collect the logit of the correct next token on both a standard forward pass and on a run where an attention head is resample ablated. The difference between the "clean" and "ablated" logit of the correct next token $\Delta\text{logit}$ captures how much the correct prediction was impacted as a result of the ablation. We also collect the head’s clean and ablated direct effect on both runs and measure the change from the clean to the new ablated direct effect, $\Delta\text{DE}_{\text{head}}$.
If the output of an attention head has no ’indirect’ downstream effects - no downstream model component depends on the output of the attention head - the change in the logit of the correct token would only be the change in the direct effect, i.e. $\Delta\text{logit}=\Delta\text{DE}_{\text{head}}$. In practice, though, self-repair occurs, and the logit difference is less in magnitude than the change in direct effect; thus, we can measure self-repair by measuring the difference between the actual logit difference and the expected amount:
$\Delta\text{logit}=\Delta\text{DE}_{\text{head}}+\text{self repair}$ (1)
@@
As resample ablating a head changes its output to one produced from an (likely) unrelated prompt, the new direct effect when resample ablating an arbitrary head is near zero ($\Delta\text{DE}_{\text{head}}\approx-DE_{\text{head}}$ when resample ablating most attention heads). As such, for convenience, we often plot the original direct effect rather than the change in direct effect. Note that if the original direct effect of an attention head was negative, we also expect the self-repair upon the head’s ablation to be negative. However, if the original direct effect is positive and the self-repair is negative, this should not be thought of as self-repair, but rather downstream breakage (where later components perform worse as a result of the ablation, such as due to the breaking of downstream circuitry).
@@ 2.2 Self-Repair Exists, but Imperfectly
We measure the self-repair of individual attention heads in Pythia-1B (Biderman et al., 2023) on 1 million tokens of The Pile (Gao et al., 2020), the dataset used to train Pythia-1B  [^1] For this experiment and others, unless stated otherwise, we use 1 Million tokens for experiments with the Pythia suite of models and GPT2-Small and GPT2-Medium, but smaller amounts for the other models.. For a given token and forward pass, we measure each head’s direct effect $DE_{head}$ and the outputted logit of the correct next token. We then individually resample ablate each attention head, measure the new logit of the same correct next token, and calculate the logit difference $\Delta\text{logit}$. Averaging these values across the whole dataset, we plot each head’s direct effect and logit difference in Figure 2 (see Appendix F for the equivalent graphs of other models).
@@ 2.3 Self-Repair is Noisy
Figure 3: We’ve selected four heads in Pythia-410M, and plotted the direct effect and logit difference when ablating the head across 5000 individual tokens in The Pile. Within a single head, these values can vary highly. The tokens between the included $y=-x$ line and the x-axis are self-repaired.
Often, the amount of self-repair is far less, or more, than the original direct effect of the head. If the magnitude of self-repair is quite low, it doesn’t quite make sense to declare that self-repair has meaningfully occurred. Additionally, sometimes the measured ’self-repair’ can be far larger than the original direct effect of the head, which means model performance improves upon the ablation of the head.
@@ 3 Nontrivial Self-Repair Due to LayerNorm
Given these observations of robust self-repair on the pretraining distribution, what might be some mechanisms behind self-repair? We originally expected for nearly all of the self-repair to occur because of a model’s attention heads and MLP layers. However, it turns out that even ’passive’ components of the model, such as normalizing factors, can induce self-repair. If we ablate a head and freeze the output of all downstream components except for the LayerNorm normalization factor, this recovers a significant logit difference relative to freezing all components including the normalization factor.
@@ 3.1 An Argument for LayerNorm Self-Repair
As argued in Elhage et al. 2021, we can simplify the LayerNorm operation such that we ’fold in’ LayerNorm projections to the weights of linear layers before and after the projection, leaving the nonlinear operation of scaling the residual stream by dividing by $S$, a normalization factor proportional to the residual stream’s norm  [^4] Precisely, the normalization factor is calculated by measuring the standard deviation of the residual stream across the residual stream dimension (not the sequence dimension). This is proportional to the norm.. These scaling factors are a non-trivial change to the forward passes of the model. We can write the logit difference from a resample ablation as the following (which we derive in Appendix B):
@@ 3.2 Empirical Findings of LayerNorm Self-Repair
On these tokens, the ratio between the clean and ablated LayerNorm scaling factors are almost always greater than one (90.66% of tokens for resample, 98.75% for mean, 100% for zero ablation).
This highlights how various forms of common ablations (zero, mean, and resample) can sometimes influence the scaling factors quite significantly. In particular, zero ablating may strongly change the norm of the residual stream. This has practical consequences for the self-repair of attention heads: Figure 5 highlights the self-repair that occurs ablating L11H10 in Pythia-160 with the different ablations. Recall that L11H10 is in the last attention layer and that Pythia-160M is a parallel-attention model: as such, the only component responsible for any self-repair is LayerNorm. Zero ablating can increase the logit difference, even when ablating the head while it has a positive direct effect  [^6] It’s not always the case that zero ablating has such an extreme effect. See Appendix D..
@@
self $\text{self repair}&=\Delta\text{logit}-\Delta\text{DE}_{\text{head}}\\ &=\sum_{h\in H}\Delta\text{DE}_{h}+\sum_{m\in M}\Delta\text{DE}_{m}+\Delta\text{DE}_{\text{LayerNorm}}\end{aligned}$ (2)
@@ 4.1 Erasure Occurs in Neurons
To provide an example, we take L10H11 in Pythia-160M, an attention head right before the final MLP layer, and run the model on a subset of The Pile, resample ablate the head, and measure the self-repair on each token. Then, we filter for the top 2% of instances in which the last MLP layer self-repairs the most. As the output of the MLP layer is the sum of the output of each neuron, we can also measure the self-repair of each neuron: as such, for each token, we isolate the individual neuron that self-repairs the most in the layer (for that instance) and plot its clean and ablated direct effect, colored by the direct effect of the entire last MLP layer, in Figure 7.
@@
Figure 7: For selected tokens on The Pile, we plot the clean and ablated direct effect of the top self-repairing neuron in the last MLP layer when resample ablating L10H11 in Pythia-160M (and include a histogram of the clean direct effects of the neurons). Each token is colored with the direct effect of the entire MLP layer on the clean run. The majority of the neurons are performing Anti-Erasure. Additionally, despite the MLP layer almost always contributing positively to the direct effect, the top self-repairing neuron typically performs Anti-Erasure and has a negative clean direct effect.
@@ 4.2 Sparse Sets of Neurons Perform Anti-Erasure
Figure 8: For four models and attention heads, we plot the frequency at which the top $N$th self-repairing neuron repaired Xof the direct effect (on the top 2% of the tokens by direct effect).
@@ 4.2.1 How Important is the Sparse Anti-Erasure?
Although a sparse number of neurons self-repair significantly, this doesn’t imply that these are the only self-repairing neurons. On a given ablation, many neurons have marginally decreased or increased direct effects. This means that while a small subset of the top self-repairing neurons may sum up to the ’total self-repair’, many other neurons may have changed direct effects that cancel each other out. In particular, when sorting, the sum of the largest fractions of self-repair may significantly exceed $1$ if they are compensated for by many neurons with reduced direct effects.
We suspect that when ablating the same attention heads, roughly 60% of the neurons in the final layer have negligible changes in direct effect and that 40% of the neurons have changes in self-repair which are small and cancel out, yet are non-negligible altogether when considering their cumulative absolute change in direct effects (Appendix C). This indicates that the full picture of self-repair in the entire MLP layer is more complex and that sparse Anti-Erasure is only one part of a larger story: it’s not the case that all the neurons stay stagnant besides a set of anti-erasure neurons. Instead, while sparse sets of anti-erasure neurons may help fuel large amounts of self-repair, there is also a significant amount of neurons that have smaller changes in their direct effect.
@@ 5 Discussion
In this section, we highlight some of the consequences of self-repair. We also outline the Iterative Inference Hypothesis, which we believe may help describe portions of self-repair.

### v2

@@ Abstract
Prior interpretability research studying narrow distributions has preliminarily identified self-repair, a phenomena where if components in large language models are ablated, later components will change their behavior to compensate. Our work builds off this past literature, demonstrating that self-repair exists on a variety of models families and sizes when ablating individual attention heads on the full training distribution. We further show that on the full training distribution self-repair is imperfect, as the original direct effect of the head is not fully restored, and noisy, since the degree of self-repair varies significantly across different prompts (sometimes overcorrecting beyond the original effect). We highlight two different mechanisms that contribute to self-repair, including changes in the final LayerNorm scaling factor and sparse sets of neurons implementing Anti-Erasure. We additionally discuss the implications of these results for interpretability practitioners and close with a more speculative discussion on the mystery of why self-repair occurs in these models at all, highlighting evidence for the Iterative Inference hypothesis in language models, a framework that predicts self-repair.
@@ 1 Introduction
We strengthen this prior work by investigating self-repair across the whole pretraining distribution, by focusing on individual attention heads, and by investigating the mechanisms behind self-repair on the whole distribution. Our key findings are:
@@
    A nontrivial fraction - possibly 30% - of the self-repairing of direct effects can be attributed to just the effect of ablations on the LayerNorm normalization factor.
@@
    MLP Erasure, a mechanism in which MLP layers ‘erase’ outputs from earlier model components and is known to explain self-repair (McGrath et al., 2023), is powered by a sparse set of Erasure neurons. The exact neurons vary between prompts.
We end with some discussion on the implications of these results for interpretability research and some present preliminary evidence for the Iterative Inference hypothesis, a speculative framework that predicts self-repair as a side effect. All the code for the experiments used in this paper is provided at https://github.com/starship006/backup_research.
@@ 2 Self-Repair on the Full Distribution Exists, but is Incomplete and Noisy
In this section, we confirm that across models, individual attention heads are ‘self-repaired’ on a general pretraining distribution. We detail how we measure self-repair, demonstrate how individual attention heads are imperfectly self-repaired across the entire pretraining distribution, and highlight how it is noisy.
2.1 Defining Self-Repair
We first detail how we define direct effect and self-repair. Our methodology aligns with prior research on self-repair, although not mirroring it identically (we detail how our work differs from prior works in Section 6). Let $(x_{1},x_{2},...,x_{n-1},x_{n})$ be a sequence of $n$ tokens we pass into a language model. A language model maps each contiguous prefix of input tokens $(x_{1},x_{2},...,x_{k-1},x_{k})$ to a final activation $r_{k}$, ultimately producing $n$ final activations $(r_{1},r_{2},...,r_{n-1},r_{n})$ for each prefix.
We follow Elhage et al. 2021 in calling these the final residual streams (as they are the output of the final MLP layer plus the final skip connection). An important fact about residual streams is that each final residual stream $r_{k}$ can be decomposed into the sum of the output of each layer, plus the original embeddings (Elhage et al., 2021). We refer to anything which adds its output into the residual stream as a model component - components can include attention layers, individual attention heads, MLP layers, or the individual neurons within them (for simplicity, we also count the model embeddings as a component). Formally, for a decomposition into components $C^{i}$, define $C^{i}_{k}$ as the output on position $k$, $r_{k}=\sum_{i}C^{i}_{k}$.
For a given final residual stream activation $r_{k}$, the model produces a distribution of logits $l_{k}$, obtained from then applying a normalization function $LN$, such as LayerNorm or RMSNorm, to the residual stream corresponding to $k$-th position, $res\in R^{d_{model}}$ (where $d_{model}$ is the dimensionality of the model’s hidden states), and multiplying it by the unembedding matrix $W_{U}\in\mathbb{R}^{V\times d_{model}}$ of the model: $l_{k}=W_{U}\cdot LN(r_{k})$. We center the logits so the average logit is zero, which prevents ‘identical translations in logits’ from occuring (see Appendix E).
$W_{U}$ is a linear map and we can take a linear approximation to $LN$ on a given input (Elhage et al., 2021). The composition of two linear functions is a linear function, i.e. a matrix multiplication plus a bias, so $l_{k}=Wr_{k}+b$ for some $W$ and $b$ (which vary with the input). Because $r_{k}$ is the sum of component outputs, $l_{k}=W(\sum_{i}C^{i}_{k})+b=(\sum_{i}WC^{i}_{k})+b$.
We often just care about the logit of the correct next token $x_{k+1}$, which corresponds to a single element of $l_{k}$. This is $logit_{clean}=W[x_{k+1}]^{T}\cdot r_{k}+b[x_{k+1}]$, where $W[x_{k+1}]$ is the $x_{k+1}$-th row of $W$ and $b[x_{k+1}]$ is the $x_{k+1}$-th element of $b$ (which are a $\mathbb{R}^{d_{model}}$ vector and a scalar).
Using the decomposition of $l_{k}$, we can define $W[x_{k+1}]^{T}\cdot C^{i}_{k}$ as the direct effect of component $i$ on position $k$. Intuitively, a component can help a model predict the final answer by either producing intermediate representations that are used by later components, or by directly boosting the correct next token: the direct effect captures the latter effect only, measuring how much a component’s output $C_{k}^{i}$ writes its output in the direction corresponding to the next correct token, $W[x_{k+1}]$.
@@ 2.1 Defining Self-Repair
Notice that just as the residual stream can be mostly decomposed into the sum of outputs of all model components, the logit of the correct next token $logit_{clean}$ can be mostly decomposed into the sum of the direct effects of all model components. Ideally, if a component has a high direct effect, it was likely ‘important’ for predicting the next token.
We contrast the direct effect with ablation-based metrics. Intuitively, ablation-based metrics measure the full effect of how a component helps the model produce the output, both via producing intermediate representations used by later components and the direct effect. We perform resample ablations, a technique that replaces the output of an head as if it were from a different set of pretraining tokens (Chan et al., 2022). We run the model again on a different set of input tokens $(x^{\prime}_{1},x^{\prime}_{2},...,x^{\prime}_{n-1},x^{\prime}_{n})$ and store the new output of an attention head $C^{\prime}_{out}$. Then, when running the model on the original set of tokens $(x_{1},x_{2},...,x_{n-1},x_{n})$, we intervene during the forward pass to replace the original output of the head $C_{out}$ with the new output $C^{\prime}_{out}$, and then continue the forward pass, a technique known as activation patching or causal mediation analysis (Vig et al., 2020).
Let the new logit distribution under this ablation be $l^{\prime}_{k}$ and the new logit of the correct next token be $logit_{ablated}$. Notice that we can also compute the ablated direct effect of the attention head. If the output of an attention head has no ‘indirect’ downstream effects—i.e., no downstream model component depends on the output of the attention head—then the change in the correct logit, $\Delta logit=logit_{ablated}-logit_{clean}$, would equal the change in the direct effect of the attention head, $\Delta DE_{head}$. However, in practice, these are rarely equal, and often $|\Delta DE_{head}|>|\Delta logit|$. This discrepancy is caused by later components changing their direct effects in a way that compensates for the resample ablated component, the phenomena we refer to as self-repair. We measure the ‘self-repair’ that occurs as:
[…]
@@
When we resample ablate a head, the new activation comes from a prompt drawn from the pre-training distribution, and so is likely to be unrelated. The new direct effect when resample ablating an arbitrary head is near zero so $\Delta\text{DE}_{\text{head}}\approx-DE_{\text{head}}$ when resample ablating most attention heads. As such, for convenience we often plot the original direct effect rather than change in direct effect. For example, consider an attention head with a direct effect of $0.5$ on the correct output token: if this head is ablated and the logit difference between the original and ablated model is only $-0.2$ logits, we can approximate the self-repair as $-0.2-(-0.5)=0.3$.
Note that our definition allows for self-repair to be negative. If the original direct effect of an ablated attention head was negative, we expect the self-repair to also be negative. However, if the original direct effect is positive and the self-repair is negative, we believe this should not be thought of as self-repair, but rather downstream breakage where later components perform worse as a result of the ablation. In this work we do not focus on studying downstream breakage.
@@ 2.2 Self-Repair Exists, but Imperfectly
We measure the self-repair of individual attention heads in Pythia-1B (Biderman et al., 2023) on 1 million tokens of The Pile (Gao et al., 2020), the dataset used to train Pythia-1B  [^1] For this experiment and others, unless stated otherwise, we use 1 Million tokens for experiments with the Pythia suite of models and GPT2-Small and GPT2-Medium, but smaller amounts for the other models.. For a given token and forward pass, we measure each head’s direct effect $DE_{head}$ and the outputted logit of the correct next token. We then individually resample ablate each attention head, measure the new logit of the same correct next token, and calculate the logit difference $\Delta\text{logit}$. Averaging these values across the whole dataset, we plot each head’s direct effect and logit difference in Figure 2 (see Appendix H for the equivalent graphs of other models).
@@ 2.3 Self-Repair is Noisy
Figure 3: We’ve handpicked four heads in Pythia-410M, and plotted the direct effect and logit difference when ablating the head across 5000 individual tokens in The Pile. Within a single head, these values can vary highly. The tokens between the included $y=-x$ line and the x-axis are self-repaired.
Often, the amount of self-repair is far less, or more, than the original direct effect of the head. If the magnitude of self-repair is quite low, it doesn’t quite make sense to declare that self-repair has meaningfully occurred. Additionally, sometimes the measured ‘self-repair’ can be far larger than the original direct effect of the head, which means model performance improves upon the ablation of the head.
@@ 3 Nontrivial Self-Repair Due to LayerNorm
Given these observations of robust self-repair on the pretraining distribution, what might be some mechanisms behind self-repair? We originally expected for nearly all of the self-repair to occur because of a model’s attention heads and MLP layers. However, it turns out that even ‘passive’ components of the model, such as normalizing factors, can induce self-repair. If we ablate a head and freeze the output of all downstream components except for the LayerNorm normalization factor, this recovers a significant logit difference relative to freezing all components including the normalization factor.
@@ 3.1 An Argument for LayerNorm Self-Repair
As argued in Elhage et al. 2021, we can simplify the LayerNorm operation such that we ‘fold in’ LayerNorm projections to the weights of linear layers before and after the projection, leaving the nonlinear operation of scaling the residual stream by dividing by $S$, a normalization factor proportional to the residual stream’s norm  [^4] Precisely, the normalization factor is calculated by measuring the standard deviation of the residual stream across the residual stream dimension (not the sequence dimension). This is proportional to the norm.. These scaling factors are a non-trivial change to the forward passes of the model. We can write the logit difference from a resample ablation as the following (which we derive in Appendix B):
@@ 3.2 Empirical Findings of LayerNorm Self-Repair
On these tokens, the ratio between the clean and ablated LayerNorm scaling factors are almost always greater than one (90.66% of tokens for resample, 98.75% for mean, 100% for zero ablation). This increase causes self-repair (even for a ratio such as 1.02, see Appendix G)!
This highlights how various forms of common ablations (zero, mean, and resample) can sometimes influence the scaling factors quite significantly. In particular, zero ablating may strongly change the norm of the residual stream. This has practical consequences for the self-repair of attention heads: Figure 5 highlights the self-repair that occurs ablating L11H10 in Pythia-160M (selected as it shows our point well) with the different ablations. Recall that L11H10 is in the last attention layer and that Pythia-160M is a parallel-attention model: as such, the only component responsible for any self-repair is LayerNorm. Zero ablating can increase the logit difference, even when ablating the head while it has a positive direct effect. Though, zero ablating doesn’t always have an extreme effect, see Appendix D.
@@
self repair $=\Delta\text{logit}-\Delta\text{DE}_{\text{head}}$ (2)
$=\sum_{h\in H}\Delta\text{DE}_{h}+\sum_{m\in M}\Delta\text{DE}_{m}+\Delta\text{DE}_{\text{LayerNorm}}$
@@ 4.1 Erasure Occurs in Neurons
A head that highlights this well is L10H11 in Pythia-160M, an attention head right before the final MLP layer. We run the model on a subset of The Pile, resample ablate the head, and measure the self-repair on each token. Then, we filter for the top 2% of instances in which the last MLP layer self-repairs the most. As the output of the MLP layer is the sum of the output of each neuron, we can also measure the self-repair of each neuron: as such, for each token, we isolate the individual neuron that self-repairs the most in the layer (for that instance) and plot its clean and ablated direct effect, colored by the direct effect of the entire last MLP layer, in Figure 7.
@@
Figure 7: On select tokens, we plot the clean and ablated direct effect of the top self-repairing neuron in the last MLP layer when resample ablating L10H11 in Pythia-160M (and include a histogram of the clean direct effects of the neurons). Each token is colored with the direct effect of the entire MLP layer on the clean run. The majority of the neurons are performing Anti-Erasure. Additionally, despite the MLP layer almost always contributing positively to the direct effect, the top self-repairing neuron typically performs Anti-Erasure and has a negative clean direct effect.
@@ 4.2 Sparse Sets of Neurons Perform Anti-Erasure
Figure 8: For four models and random attention heads, we plot the frequency at which the top $N$th self-repairing neuron repaired Xof the direct effect (on the top 2% of the tokens by direct effect).
@@ 4.2.1 How Important is the Sparse Anti-Erasure?
Although a sparse number of neurons self-repair significantly, this doesn’t imply that these are the only self-repairing neurons. On a given ablation, many neurons have marginally decreased or increased direct effects. This means that while a small subset of the top self-repairing neurons may sum up to the ‘total self-repair’, many other neurons may have changed direct effects that cancel each other out. In particular, when sorting, the sum of the largest fractions of self-repair may significantly exceed $1$ if they are compensated for by many neurons with reduced direct effects.
We suspect that when ablating the same attention heads, roughly 60% of the neurons in the final layer have negligible changes in direct effect and that 40% of the neurons have changes in self-repair which are small and cancel out, yet are non-negligible when considering their combined absolute change in direct effects (Appendix C). This indicates that the full picture of self-repair in the entire MLP layer is more complex and that sparse Anti-Erasure is only one part of a larger story: it’s not the case that all the neurons stay stagnant besides a set of anti-erasure neurons. Instead, while sparse sets of anti-erasure neurons may help fuel large amounts of self-repair, there is also a significant amount of neurons that have smaller changes in their direct effect.
@@ 5 Discussion
In this section, we discuss some of the consequences of self-repair. We also outline the Iterative Inference Hypothesis, which we believe may help describe portions of self-repair.